OpenStudy (anonymous):

Factor completely 50a*2b*5 − 35a*4b*3 + 5a*3b*4

OpenStudy (mathmath333):

take 5 common \(\large \color{black}{\begin{align}& 50a^{2}b^{5} − 35a^{4}b^{3} + 5a^{3}b^{4}\hspace{.33em}\\~\\ &= 5(10a^{2}b^{5} − 7a^{4}b^{3} + a^{3}b^{4})\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (mathmath333):

this can also be written as \(\large \color{black}{\begin{align} &= 5(10a^{2}b^{5} − 7a^{2}\cdot a^{2} b^{3} + a^{2}\cdot a^{1}b^{4})\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (mathmath333):

u can take a^2 common now \(\large \color{black}{\begin{align} &= 5a^{2}(10b^{5} − 7 a^{2} b^{3} + ab^{4})\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (mathmath333):

try to take \(b\) common now

OpenStudy (anonymous):

im so confused

OpenStudy (mathmath333):

which is the smallest of \(b^{4}\) ,\(b^{3}\) and \(b^{5}\)

OpenStudy (mathmath333):

or which one is small, 4,3 or 5

OpenStudy (mathmath333):

is 3 smallest

OpenStudy (mathmath333):

??